# Shadows of Rotating Black Hole Solution with a Dark Matter Halo

### v1

@@ Abstract
In this paper, we start from static black hole (BH) immersed in a Dehnen-type dark matter (DM) halo and employ the Newman-Janis algorithm (NJA) to generate the rotating black hole solution with a dark matter halo. Then we study optical properties of newly obtained rotating BH in DM halo, including the shadow’s geometrical shape, photon sphere and the dependence of the shadow radius from DM parameters. Additionally, assuming that spacetime of a supermassive black hole (SMBH) is described by the newly obtained rotating BH solution we analyze the parameters of the model with shadow size estimates based on the Event Horizon Telescope (EHT) observations of M87 and SgrA SMBHs.
I Introduction
@@
One of the most fascinating predictions of General Relativity (GR) was the formation of black holes (BH) at the end state of evolution of massive stars and the assumption of existence of black holes in the center of galaxies Abbott (2017); Abbott and Abbott (2016) which are responsible for high energetics of active galactic nuclei (AGN). The last decade was related to the triumphal revolutionary discoveries in relativistic astrophysics. However, these essential observations were not enough to explain the nature of the interaction between black holes and their surroundings. Therefore, understanding a behavior of the dark matter (DM) is crucial task in GR as there are enough evidence that BHs are surrounded by DM halo Iocco et al. (2015); Bertone and Tait (2018). Although, numerous efforts have been done to understand the nature of the DM halo, there is still only one way, gravitational interactions, to understand DM. Thus, finding exact analytical model for DM halo is utmost problem in GR. There are several analytical models, Einasto Graham et al. (2006), Navarro-Frenk-White Navarro et al. (1996), Burkert Burkert (1995) and the Dehnen model Dehnen (1993); Shukirgaliyev et al. (2021); Al-Badawi et al. (2025); Maeda et al. (2025); Shen et al. (2025, 2024) to describe the interaction between BH and DM. In this work, we have chosen core model of the Dehnen-type dark matter halo with the known exact analytical solution for static BH in DM halo and then using Newman-Janis algorithm we generated the novel rotating black hole solution with a dark matter halo.
@@
Recent studies within the Dehnen-type DM halo framework have also examined interactions between DM and BHs from various viewpoints. In particular, these studies thoroughly examine black holes embedded within Dehnen-type dark matter halos exploring the influence of the DM halo on quasinormal modes, the BH shadow, and the photon sphere radius Xamidov et al. (2025), alongside analyzing gravitational waveforms generated by periodic orbits Alloqulov et al. (2025) for the BH-Dehnen-type DM halo solution. Furthermore, the null geodesics and the thermodynamics of the effective BH-DM halo system wer studied in Gohain et al. (2024), which is followed by an analysis that constrains the parameters of the DM halo Xamidov et al. (2025).
@@
Observations from the EHT, specifically images of supermassive black holes (SMBHs) Sgr A and M87, give opportunities to compare our theoretical findings with the real observational data Vagnozzi and Visinelli (2019). In 2019, the EHT collaboration made a groundbreaking announcement by providing the first image of a black hole, a shadow-like representation of the supermassive black hole located in the center of the M87 galaxyAkiyama and et al. (2019) (Event Horizon Telescope Collaboration). Following, researchers from the EHT project unveiled an image of the black hole at the center of the Milky Way galaxy, known as Sgr A* in 2022. In accordance with GR, the observed images of the two black holes, M87 and Sgr A, are consistent with the properties predicted for a Kerr black hole. This agreement indicates that these black holes are effectively characterized by the Kerr metric, which serves as the solution to the Einstein field equations for a rotating black holeAkiyama and et al. (2022) (Event Horizon Telescope Collaboration). Although Kerr-like black holes emerging from modified gravity theories or including a DM halo are not fully validated due to the relative deviation in quadrupole moments and the current measurement uncertainties in spin or angular momentum, they cannot be completely excluded Cardoso and Pani (2019). In recent years, the study of black hole shadows has attracted significant attention, particularly in the context of modified gravity theories Zahid et al. (2022). Additionally, the exploration of spinning black holes has offered new insights into the optical properties and shadow features of these celestial objects Zubair et al. (2022). Moreover, observational data, including those gathered by the EHT, have been crucial in constraining the range of theoretical parametersRaza et al. (2024). Consequently, one of the key challenges in modern astrophysics is evaluating the nature and existence of dark matter models using the data supplied by the EHT Collaboration. In this paper, we obtain constraints for the characteristic density $\rho_{s}$ and the characteristic scale factor $r_{s}$ of the DM halo using the observed data mentioned above.
@@
The paper organized as follows: in Sec. II, we generate a solution for a rotating BH surrounded by a Dehnen-type dark matter halo using Newman-Janis algorithm. In addition, event horizon of the rotating BH with a dark matter halo is analyzed in this section. The following null geodesics equations are derived using the Hamilton-Jacobi equation, the photon sphere around a rotating BH in a dark matter halo is analyzed, and finally, the shadow of the rotating BH in dark matter halo is investigated in Sec. III. Then we obtain constraints for the BH parameters $r_{s}$ and $\rho_{s}$ applying the EHT observation data to the obtained theoretical results in Sec. IV. Finally, we give our main conclusions in Sec. V.
@@
II Spacetime of rotating black hole with the dark matter halo
@@
The authors in Uktamov et al. (2025) have found a novel static black hole solution surrounded by Dehnen-type dark matter halo in the form:
@@
$f(r)=1-\frac{2M}{r}-\frac{8\pi\rho_{s}r_{s}^{2}}{1+\frac{r_{s}}{r}}\left(1+\frac{r_{s}}{r}\right)\log{\left(1+\frac{r_{s}}{r}\right)}\ ,$ (2)
@@
Now using Newman and Janis method (NJA) Toshmatov et al. (2017); Newman and Janis (1965) we will generate rotating BH space-time in dark matter halo. To do it, we start with coordinate transformation from the Boyer-Lindquist (BL) coordinates $(t,r,\theta,\phi)$ to the Eddington-Finkelstein (EF) coordinates $(u,r,\theta,\phi)$:
@@
III Shadow of rotating black hole in Dehnen-type DM halo
@@
where we have introduced new variables $\eta=\frac{\mathcal{K}}{E^{2}}$ and $\xi=\frac{L}{E}$. Then we can use condition for unstable circular orbits $V_{eff}(r_{ph})=\frac{\partial V_{eff}}{\partial r}\mid_{r_{ph}}=0$, where $r_{ph}$ is the radius of the photon sphere, to find $\eta$ and $\xi$ as:
@@
Now, to obtain the silhouette of the BH shadow, one may plot the line described by $\xi$ vs. $\eta$ in celestial coordinates.
@@
IV Constraints from EHT observational data
Now we will obtain constraints from the observational data of the EHT on the image characteristics of SMBHs. To do it, we ought to find a closed area of the shadow silhouette as Zahid et al. (2023) (Abdujabbarov et al. (2015)):

### v2

# New analytical model of rotating black hole with dark matter halo: constraints from EHT observations and accretion disk

@@ Abstract
In this paper, we start from a static black hole (BH) immersed in a Dehnen-type dark matter (DM) halo and employ the Newman-Janis algorithm (NJA) to generate the rotating black hole solution with a dark matter halo. Also, we have checked the validity of the obtained space-time. Then we study optical properties of newly obtained rotating BH in DM halo, including the shadow’s geometrical shape, deflection angle of light based Ono, Ishihara and Asada (OID) method, photon sphere and the dependence of the shadow radius on DM parameters. Additionally, assuming that spacetime of a supermassive black hole (SMBH) is described by the newly obtained rotating BH solution, we analyze the parameters of the model with shadow size estimates based on the Event Horizon Telescope (EHT) and Gravity collaboration observations of M87 and Sgr A SMBHs. Then we have used Markov Chain Monte Carlo (MCMC) analysis to constrain DM parameters $\rho_{s}$, $r_{s}$ and BH mass $M$, BH spin $a$, also we show that best-fit values for the parameters $\rho_{s}$, $r_{s}$ are well agreement with previous results which indicate physically reasonability of the our model. Finally, we have analyzed the electromagnetic radiation flux of the rotating BH in the DM halo employing a ray tracing code.
keywords:
Dehnen-type dark matter halo , Newman-Janis algorithm , shadow , thin accretion disk
journal: Journal Name
1 Introduction
@@
One of the most fascinating predictions of General Relativity (GR) was the formation of black holes (BH) at the end state of evolution of massive stars and the assumption of existence of black holes in the center of galaxies Abbott et al. [2016], Abbott and Abbott [2016] which are responsible for high energetics of active galactic nuclei (AGN). The last decade was related to the triumphal revolutionary discoveries in relativistic astrophysics. However, these essential observations were not enough to explain the nature of the interaction between black holes and their surroundings. Therefore, understanding a behavior of the dark matter (DM) is crucial task in GR as there are enough evidence that BHs are surrounded by DM halo Iocco et al. [2015], Bertone and Tait [2018]. Although, numerous efforts have been done to understand the nature of the DM halo, there is still only one way, gravitational interactions, to understand DM. Thus, finding exact analytical model for DM halo is utmost problem in GR. There are several analytical models, Einasto Graham et al. [2006], Navarro-Frenk-White Navarro et al. [1996], Burkert Burkert [1995] and the Dehnen model Dehnen [1993], Shukirgaliyev et al. [2021], Al-Badawi et al. [2025], Maeda et al. [2025], Shen et al. [2025, 2024] to describe the interaction between BH and DM. In this work, we have chosen core model of the Dehnen-type dark matter halo with the known exact analytical solution for static BH in DM halo and then using Newman-Janis algorithm we generated the novel rotating black hole solution with a dark matter halo.
@@
Recent studies within the Dehnen-type DM halo framework have also examined interactions between DM and BHs from various viewpoints. In particular, these studies thoroughly examine black holes embedded within Dehnen-type dark matter halos exploring the influence of the DM halo on quasinormal modes, the BH shadow, and the photon sphere radius Xamidov et al. [2025], alongside analyzing gravitational waveforms generated by periodic orbits Alloqulov et al. [2025] for the BH-Dehnen-type DM halo solution. Furthermore, the null geodesics and the thermodynamics of the effective BH-DM halo system wer studied in Gohain et al. [2024], which is followed by an analysis that constrains the parameters of the DM halo Xamidov et al. [2025].
@@
Observations from the EHT, specifically images of supermassive black holes (SMBHs) Sgr A and M87, give opportunities to compare our theoretical findings with the real observational data Vagnozzi and Visinelli [2019], Bambi et al. [2019], Vagnozzi et al. [2023]. In 2019, the EHT collaboration made a groundbreaking announcement by providing the first image of a black hole, a shadow-like representation of the supermassive black hole located in the center of the M87 galaxyAkiyama and et al. (2019) [Event Horizon Telescope Collaboration]. Following, researchers from the EHT project unveiled an image of the black hole at the center of the Milky Way galaxy, known as Sgr A*, in 2022. In accordance with GR, the observed images of the two black holes, M87 and Sgr A, are consistent with the properties predicted for a Kerr black hole. This agreement indicates that these black holes are effectively characterized by the Kerr metric, which serves as the solution to the Einstein field equations for a rotating black holeAkiyama and et al. (2022) [Event Horizon Telescope Collaboration]. Although Kerr-like black holes emerging from modified gravity theories or including a DM halo are not fully validated due to the relative deviation in quadrupole moments and the current measurement uncertainties in spin or angular momentum, they cannot be completely excluded Cardoso and Pani [2019]. In recent years, the study of black hole shadows has attracted significant attention, particularly in the context of modified gravity theories Zahid et al. [2022]. Additionally, the exploration of spinning black holes has offered new insights into the optical properties and shadow features of these celestial objects Zubair et al. [2022]. Moreover, observational data, including those gathered by the EHT, have been crucial in constraining the range of theoretical parametersRaza et al. [2024]. Consequently, one of the key challenges in modern astrophysics is evaluating the nature and existence of dark matter models using the data supplied by the EHT Collaboration. In this paper, we obtain constraints for the characteristic density $\rho_{s}$ and the characteristic scale factor $r_{s}$ of the DM halo using the observed data mentioned above.
@@
The paper organized as follows: in Sec. 2, we generate a solution for a rotating BH surrounded by a Dehnen-type dark matter halo using Newman-Janis algorithm. In addition, event horizon of the rotating BH with a dark matter halo is analyzed in this section. The following null geodesics equations are derived using the Hamilton-Jacobi equation, the photon sphere around a rotating BH in a dark matter halo is analyzed, and finally, the shadow of the rotating BH in dark matter halo is investigated in Sec. 3. Subsequently, deflection angle of light by a rotating BH in DM halo and effect of the DM halo on deflection angle $\hat{\alpha}$ is investigated in Sec.4. Then we obtain constraints for the BH parameters $r_{s}$ and $\rho_{s}$ applying the EHT and Gravity collaboration observation data to the obtained theoretical results in Sec. 5 and then we employed MCMC analysis to find best-fit values of the DM parameters $\rho_{s}$, $r_{s}$. Also we obtain image of the thin accretion in rotating BH surrounded by Dehnen-type DM halo in Sec.  6. Finally, we give our main conclusions in Sec. 7.
@@
2 Spacetime of rotating black hole with the dark matter halo
@@
The authors in Uktamov et al. [2025a] have found a novel static black hole solution surrounded by Dehnen-type dark matter halo in the form:
@@
$f(r)=1-\frac{2M}{r}-8\pi\rho_{s}r_{s}^{2}\log{\left(1+\frac{r_{s}}{r}\right)}\ ,$ (2)
@@
Now using Newman and Janis method (NJA) Toshmatov et al. [2017], Azreg-Aïnou [2014, 2014, 2014], Newman and Janis [1965] we will generate rotating BH space-time in dark matter halo. To do it, we start with coordinate transformation from the Boyer-Lindquist (BL) coordinates $(t,r,\theta,\phi)$ to the Eddington-Finkelstein (EF) coordinates $(u,r,\theta,\phi)$:
@@
2.1 Checking validity of the obtained rotating black hole solution
Now our aim is to check whether our obtained rotating black hole solution with dark matter halo satisfies the Einstein field equations or not. Using the line element of the black hole surrounded by Dehnen-type DM halo (8,9), the components of the Einstein tensor can be expressed as:
where we have introduced a new variable as $F(r)=\frac{r^{2}\left[1-f(r)\right]}{2}$. The Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$ could then be checked using an appropriate form of orthogonal bases (Azreg-Aïnou [2014], Jusufi et al. [2019]) such as:
$e_{t}^{\mu}=-\frac{\left(r^{2}+a^{2},0,0,a\right)}{\sqrt{\Sigma\Delta}}\,,\,\,e_{r}^{\mu}=-\frac{\sqrt{\Delta}\left(0,1,0,0\right)}{\sqrt{\Sigma}}\,$, (11a)
$e_{\theta}^{\mu}=-\frac{\left(0,0,1,0\right)}{\sqrt{\Sigma}}\,,\,\,e_{\phi}^{\mu}=\frac{\left(a\sin^{2}{\theta},0,0,1\right)}{\sqrt{\Sigma}\sin{\theta}}\,.$ (11b)
Also, according to (Burinskii et al. [2002], Azreg-Aïnou [2014]) the components of the energy-momentum tensor $T_{\mu}^{\nu}=\text{diag}\left[-\rho,p_{r},p_{\theta},p_{\phi}\right]$ can be expressed as:
$-\rho=p_{r}=\frac{\left(rF_{,r}-F\right)}{4\pi\Sigma^{2}}\,,\,\,p_{\theta}=p_{\phi}=-p_{r}+\frac{F_{,rr}}{8\pi\Sigma}.$ (12a)
Then one can easily notice that the newly obtained rotating black hole solution with a DM halo (8) satisfies the Einstein field equations indicating that the obtained line element is physically valid as $\rho=\frac{1}{8\pi}e_{t}^{\mu}e_{t}^{\nu}G_{\mu\nu}$, $p_{r}=\frac{1}{8\pi}e_{r}^{\mu}e_{r}^{\nu}G_{\mu\nu}=\frac{1}{8\pi}g^{rr}G_{rr}$, $p_{\theta}=\frac{1}{8\pi}e_{\theta}^{\mu}e_{\theta}^{\nu}G_{\mu\nu}=\frac{1}{8\pi}g^{\theta\theta}G_{\theta\theta}$, $p_{\phi}=\frac{1}{8\pi}e_{\phi}^{\mu}e_{\phi}^{\nu}G_{\mu\nu}$.
3 Shadow of rotating black hole in Dehnen-type DM halo
@@
where we have introduced new variables $\eta=\frac{\mathcal{K}}{E^{2}}$ and $\xi=\frac{L}{E}$. Then we can use condition for unstable circular orbits $V_{eff}(r_{ph})=\frac{\partial V_{eff}}{\partial r}\mid_{r_{ph}}=0$, where $r_{ph}$ Khodadi et al. [2024] is the radius of the photon sphere, to find $\eta$ and $\xi$ as:
@@
Now, to obtain the silhouette of the BH shadow, one may plot the line described by $\xi$ vs. $\eta$ in celestial coordinates Tsukamoto et al. [2014].
@@
4 Deflection angle of light by a rotating black hole in Dehnen-type DM halo.
In this section, we will study the deflection angle for a rotating black hole in Dehnen-type DM halo via OIA method (Ono et al. [2017], Övgün et al. [2019]). To achieve this, we begin by applying the null condition $ds^{2}=0$ and derive the orbital equation in the equatorial plane, where $\theta=\frac{\pi}{2}$. We also utilize a bounded, two-dimensional orientable surface, as depicted in Fig. 5. Following, the geodesic constants, energy $E=-k_{(t)}^{\mu}p_{\mu}$ and angular momentum $L=k_{(\phi)}^{\mu}p_{\mu}$in which:
$k_{t}^{\mu}=(1,0,0,0)\,\,\,\,\text{timelike Killing vector}\,$,
$k_{\phi}^{\mu}=(0,0,0,1)\,\,\,\,\text{spacelike Killing vector}\,$,
for the line element (8) can be expressed as:
$E=-g_{tt}\dot{t}-g_{t\phi}\dot{\phi}\,$, (23a)
$L=g_{t\phi}\dot{t}+g_{\phi\phi}\dot{\phi}\,$, (23b)
here dot means derivative with respect to affine parameter.
Figure 5: The geometry of a quadrilateral embedded in curved space is presented.Övgün et al. [2019]
Subsequently, impact parameter $b$ can be defined as:
$b=\frac{L}{E}=\frac{g_{t\phi}+g_{\phi\phi}\frac{d\phi}{dt}}{-g_{tt}-g_{t\phi}\frac{d\phi}{dt}}\,$, (24)
which paves a way to find $\frac{dr}{d\phi}$ using condition $ds^{2}=0$ for the line element (8) as:
$(\frac{dr}{d\phi})^{2}=\frac{(g_{t\phi}^{2}-g_{tt}g_{\phi\phi})(b^{2}g_{tt}+2bg_{t\phi}+g_{\phi\phi})}{g_{rr}(g_{tt}b+g_{t\phi})^{2}}\,$, (25)
or after introducing new variable as $u=\frac{1}{r}$ Eq. (25) can be rewritten as:
$(\frac{du}{d\phi})^{2}=\frac{u^{4}(g_{t\phi}^{2}-g_{tt}g_{\phi\phi})(b^{2}g_{tt}+2bg_{t\phi}+g_{\phi\phi})}{g_{rr}(g_{tt}b+g_{t\phi})^{2}}\,.$ (26)
Again, using the null condition $ds^{2}=0$ for metric (8) we are able to find $dt$ as:
$dt=\sqrt{\gamma_{ij}dx^{i}dx^{j}}+\beta_{i}dx^{i}\,$, (27a)
$dl^{2}=\gamma_{ij}dx^{i}dx^{j}=\frac{g_{t\phi}^{2}-g_{\phi\phi}g_{tt}}{g_{tt}^{2}}d\phi^{2}-\frac{g_{rr}}{g_{tt}}dr^{2}\,$, (27b)
$\beta_{i}dx^{i}=-\frac{g_{t\phi}}{g_{tt}}d\phi\,$, (27c)
in which $\gamma_{ij}\neq g_{ij}\,\,(i,j=1,2,3)$ is the spatial metric and $l$ is the arc-length.
Then, on the equatorial plane the photon’s path is described by a unit tangent vector $e^{i}$, given by(Övgün et al. [2019],Ono et al. [2017]):
$e^{i}=\frac{1}{\chi}\left(\frac{dr}{d\phi},0,1\right)\,$, (28)
in which $\chi$ can be obtained using condition $\gamma e^{i}e^{j}=1$ as
$\frac{1}{\chi}=\frac{g_{tt}(g_{tt}b+g_{t\phi})}{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}\,.$ (29)
In the case where the unit radial vector is directed outgoing, we find:
$R^{i}=\left(\frac{1}{\sqrt{\gamma_{rr}},},0\,,0\,\right)\,$, (30)
which enables to find angle computed from the outgoing radial direction as:
$\sin{\Psi}=-\frac{g_{t\phi}+bg_{tt}}{\sqrt{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}}\,$, (31)
where we have used expression $\cos{\Psi}=\gamma_{ij}e^{i}R^{j}$. Eq.(31) can be expressed for metric (9) in the form:
$\sin{\Psi}=\frac{b}{r}\times\frac{1-\frac{2\tilde{M}}{r}+\frac{2a\tilde{M}}{br}}{\sqrt{1-\frac{2\tilde{M}}{r}+\frac{a^{2}}{r^{2}}}}\,$, (32)
here $\tilde{M}=\left[M+M_{D}\log{(1+\frac{r_{s}}{r})^{1+\frac{r}{r_{s}}}}\right]$, eq.(32) converts to the exactly Kerr form (Ono et al. [2017]) when $\rho_{s}\to 0$. Subsequently, eq.(32) can be aproximated as:
$\sin{\Psi}=\frac{b}{r}\left(1-\frac{\tilde{M}}{r}+\frac{2a\tilde{M}}{br}\right)+\mathcal{O}(\frac{\tilde{M}^{2}}{r^{2}},\frac{a^{2}}{r^{2}})\,.$ (33)
Then we are able to obtain relation between the angles at the receiver $\Psi_{R}$ and the source $\Psi_{S}$ in the form:
$\Psi_{R}-\Psi_{S}$ $=$ $\arcsin{(bu_{R})}+\arcsin{(bu_{S})}-\pi$-
- $\frac{\tilde{M}bu_{R}^{2}}{\sqrt{1-b^{2}u_{R}^{2}}}-\frac{\tilde{M}bu_{S}^{2}}{\sqrt{1-b^{2}u_{S}^{2}}}+$
$+$ $\frac{2\tilde{M}u_{R}^{2}}{\sqrt{1-b^{2}u_{R}^{2}}}+\frac{2\tilde{M}u_{S}^{2}}{\sqrt{1-b^{2}u_{S}^{2}}}\,.$
The deflection angle can now be formulated in terms of the newly defined angles at the receiver $\Psi_{R}$, the source $\Psi_{S}$, and the coordinate angle $\phi_{RS}$ as:
$\hat{\alpha}=\Psi_{R}-\Psi_{S}+\phi_{RS}\,$, (35)
where $\phi_{RS}$ can be calculated employing Eq.(25) as:
$\phi_{RS}$ $=$ $\pi-\arcsin{(bu_{R})}-\arcsin{(bu_{S})}+$
$+$ $\frac{2\tilde{M}^{*}}{b}\left[\frac{2-b^{2}u_{R}^{2}}{2\sqrt{1-b^{2}u_{R}^{2}}}+\frac{2-b^{2}u_{S}^{2}}{2\sqrt{1-b^{2}u_{S}^{2}}}\right]$
- $\frac{2a\tilde{M}^{*}}{b^{2}}\left[\frac{1}{\sqrt{1-b^{2}u^{2}_{S}}}+\frac{1}{\sqrt{1-b^{2}u^{2}_{S}}}\right]\,$,
here we used weak field and slow rotation approximations (see, (Ono et al. [2017]), Ishihara et al. [2016] for details).
Finally, taking far limit $r_{S}\to 0,\,\,\,r_{R}\to 0$ and substituting Eqs.(4) and (4) into Eq.(35) we will have expression for deflection angle:
$\hat{\alpha}=\frac{4\tilde{M}^{*}}{b}\pm\frac{4a\tilde{M}^{*}}{b^{2}}\,$, (37)
where +,+,\,\,\,- signs here correspond to the retrograde and prograde photon orbits and $\tilde{M}^{*}\approx M+4\pi\rho_{s}r_{s}^{3}$.
Figure 6: Dependence of the deflection angle $\hat{\alpha}$ on the impact parameter $b$ for various cases.
Then we have shown how deflection angle $\hat{\alpha}$ depend on the impact parameter $b$ in Fig.(6). One can see from graphics in Fig.(6) that presence of the DM halo increases the value of the deflection angle $\hat{\alpha}$.
5 Constraints from EHT observational data
Now we will obtain constraints from the observational data of the EHT on the image characteristics of SMBHs. To do it, we ought to find a closed area of the shadow silhouette as Zahid et al. [2023] (Abdujabbarov et al. [2015]):